the length of a rectangle garden is 3ft less than 2 times it's width. If the perimeter of the garden is 54ft, find the dimension of a garden.
Answer by Paul(988)

Length = 2x-3
width = x
Perimeter = 54

SInce the formula for P is 2(l+w)=P

We plug that in.
2(2x-3+x)=54
6x-6=54
6x=60
x=10

2(10)-3=17

Hence the width is 10cm and the length is 17cm.
Paul.